# Knowledge Graph Supported Benchmark and Video Captioning for Basketball

### 25 Jan 2024 v1

@@ 5.2 Main Results
Table 4: Ablation study on MbgVC. Ki and Kn denote the player images knowledge and player names knowledge, respectively. Attn. is the multi-head self-attention module. Bi-GRU is bi-directional GRU module. Numbers in bold indicate the best performance.
5.3 Ablation Study
@@
To verify the contribution of the external knowledge and other modules in EAC model, we also perform an ablation study by progressively adding these as input. The results of ablation study are shown in Tab. 4. Model ① consists of Resnet-18 and T5 decoder, with its input being solely video. From the comparison results between of model ① and model ②, it can be observed that adding players’ images to model based on ① does not improve the model’s performance. During the training stage, if the model has not been exposed to enough specific names, or has not learned how to infer names from video and image features, it may not be able to generate these names. The comparison results between model ① and model ③ highlight a significant improvement by solely adding players’ names to the model based on ①. Providing a list of names as input allows the model to effectively integrate this information with the video features, leading to more accurate generation of names. This is likely providing the model with external contextual information to help it make more accurate predictions. From the comparison results between model ③ and model ④, we can find that player images can emphasize the roles based on entity names and improve the names’ prediction. From the comparison results of models ④-⑦, it can be observed that adding multi-head self-attention and Bi-GRU modules can improve the performance of the model. The multi-head self-attention module can model the associations among players and focus on key players. The Bi-GRU module can model the temporal information of video frame features, capturing the action information. From the above results, the external players knowledge brings significant gains in entity awareness. The model can generate text descriptions with specific player names and fine-grained actions through external knowledge and temporal modeling.
5.4 Error Analysis
The error analysis are conducted by presenting cases of EAC. As shown in Fig. 5, we show several errors. (1) Player Mismatching: when the number of players in the sentence is more than one, the model may decode the name of one of the players incorrectly. (2) Action Confusion: model tends to confuse similar-looking actions, such as layups and close-range two-point jump shots. (3) Lack Distance-aware Perception: model tends to confuse three-point jump shots and long-range two-point jump shots. For example, the player appears to be shooting a three-point shot, but is actually inside the three-point line.
@@ 5.4 Error Analysis
Given the above result, we discuss some potential ways for developing an advanced models for KABVC task. First, beyond the current object detection, it is necessary to enhance the model’s ability to understand and handle names. Second, regarding action confusion, this may be because the model is not sensitive enough to subtle differences in the video to accurately distinguish similar actions. We need to improve the feature extraction abilities of the model to more accurately recognize similar-looking actions. Third, for the lack of distance perception, this may be because the model has trouble processing spatial information, especially when judging the relative distance of players from the basket. This indicates that we need to further improve the spatial awareness of our model.
6 Conclusion
@@
In this paper, we introduce the task of knowledge aided basketball video captioning (KABVC) for live text broadcast. To achieve this task, we further construct a multimodal basketball game knowledge graph (MbgKG) to provide the external knowledge. Through partial relationships in the knowledge graph, a new multimodal basketball game video captioning dataset is then constructed. We also propose a new metric GDS for video captioning, which consider the accuracy of entity name prediction. Furthermore, our experiments validate the importance of additional knowledge, including player images and entity names, for basketball video captioning.
Appendix A Construction of MbgKG
Figure 6: An example of the play-by-play page.
Figure 7: Examples of 9 types of basketball events. Each type has the corresponding keyword marked in red.
Figure 8: Examples of player information and team information.
Figure 9: An example of the multi-tuples parsed from event descriptions, players’ information, and teams’ information. In this multi-tuples, the home team information is marked in red, the away team information is marked in blue, and events are marked in green.
The example of play-by-play page is shown in Fig. 6. This page not only contains the event descriptions, but also contains links to player information and team information. We can collect relevant players’ information and teams’ information through these links. All the collected event descriptions can be divided into 9 categories, as shown in Fig. 7. Each event has its own unique keywords which are marked in red. In addition to keywords, there are also some special words in different events. For example, in shooting events, there are words such as “assist by” and “block by”. By obtaining the indices of these keywords and special words, we can parse the descriptions into multi-tuples. Fig. 8 shows the players’ information and teams’ information. We only collect information that is available for most players, including “name”, “position”, “shoots”, “height”, “weight”, “birth”, “NBA debut”, and the link to their profile pictures. For teams’ information, we only collect “name”, “season”, “coach”, “executive”, “arena”, and the link to teams’ logos. By using OCR technology to recognize the time on the scoreboard in each video frame, we can compare it to the time on the play-by-play page. If the recognized time matches the time on the scoreboard, the video timestamp of the current event is recorded in the multi-tuples. Fig. 9 shows the example of the multi-tuples parsed from event descriptions, players’ information, and teams’ information. Through these efforts, a Multimodal Basketball Game Knowledge Graph (MbgKG) containing 11489 events and 42870 relationships is constructed. MbgKG not only contains the static information, but also includes the temporal dynamic information. We believe that this knowledge graph will continue to be extended and be explored more valuable knowledge for sports video understanding.
Appendix B Construction of MbgVC
In basketball, shot and rebound events are the most prevalent. Meanwhile, we count the number of different types of events in MbgKG. Tab. 5 shows that there are 4400 shot events and 2649 rebound events, which are the most common types of events. Based on the aforementioned common knowledge and statistical data, videos and event descriptions about shot and rebound are extracted from MbgKG to construct the dataset.
Shots can be divided into three types: two-point jump shot (2p jump shot), two-point layup (2pt-layup), and three-point jump shot (3p jump shot). Rebounds can be divided into two types: defensive rebound and offensive rebound. In addition, scoring and missing a shot are represented by “make” and “miss”, respectively. When a shot is not scored, it is often followed by a rebound event. Therefore, we combine the missed shot event and rebound event into one event. In total, there are nine types of shooting events, which are “2p-succ.”, “2p-fail.-off.”, “2p-fail.-def.”, “2p-layup-succ.”, “2p-layup-fail.-off.”, “2p-layup-fail.-def.”, “3p-succ.”, “3p-fail.-off.” and “3p-fail.-def.” Subsequently, based on these nine types of shooting events and the relationships within MbgKG, we extract videos, text descriptions, and player information from MbgKG. It’s worth noting that there are some issues with the text descriptions in MbgKG: 1) The names of the players are abbreviated, which does not meet the requirements of actual applications; 2) The texts contain descriptions of shooting distances, such as ”from xx ft” and ”at rim”. This makes it difficult for the captioning model to generate accurate distances; 3) The text descriptions do not conform to the structural and compositional rules of natural language.
Figure 10: Qualitative results on MbgVC dataset. (V2C: Video2Commonsense; C4C: Clip4Caption; SwB: SwinBert; EAC: our proposed model; GT: the ground truth). The different specific entity names are marked in red and blue, respectively. And the fine-grained actions are marked green. Since V2C does not have its own tokenizer and vocabulary list, it cannot decode names. So names are replaced with the special token $<$UNK$>$.
In response to the above issues, we design some modification rules to address them. For names, we replace abbreviated names in the text descriptions with the full names from the player attribute information in MbgKG (“L. James” $\Longrightarrow$ “LeBron James”). The distance descriptions (“from XX ft” and “at rim”) in the texts are directly removed. Sometimes, when a player scores a shot, it is assisted by his teammate. We modify “(assist by SOMEONE)” to “with an assist from SOMEONE”. When a player doesn’t score a shot, sometimes it’s because the ball is blocked by the opposing player. We modify “(block by SOMEONE)” to “blocked by SOMEONE”.
num. 986 2649 68 271 1103 1189 774 49 4400
Table 5: Statistics of the events in MbgKG. “num.” is the abbreviation of “number”. Numbers in bold indicate the highest two values.
Appendix C Evaluation Metric
Our task is oriented towards basketball live text broadcast, where concise and succinct text descriptions can allow users to quickly grasp and understand the content of the game in a short time. The names of the players and their corresponding fine-grained actions are the main components of the text descriptions. Existing evaluation metrics based on n-gram such as CIDEr, METEOR, Rouge-L, and BLEU, primarily penalize the semantically correct but textually inconsistent text descriptions. Each evaluation metric has its unique calculation method and principle, and no single metric can be considered as authoritative and comprehensive. Therefore, the most direct approach is to evaluate the quality of the text description by combining multiple metrics.
Since existing evaluation metrics cannot assess the accuracy of player name prediction, we introduce $e_{acc}$ to address this issue. When evaluating concise texts, the effectiveness of Rouge-L diminishes as it relies on the length of the longest common subsequence between the two texts. Therefore, compared to other evaluation metrics, the importance of Rouge-L would be somewhat lessened. We believe that the assessment of player name prediction is crucial, as it is a fundamental requirement in real-world applications. Based on the aforementioned analysis, we assign the highest coefficient to $e_{acc}$ and the lowest coefficient to Rouge-L. The coefficients of CIDEr and METEOR fall between those of $e_{acc}$ and Rouge-L. From experience, the coefficient of eacc is set to 4, the coefficients of CIDEr and METEOR are set to 2, and the coefficient of Rouge-L is set to 1. We sought opinions from several experts and scholars in the field of video captioning, all of whom expressed that the coefficient allocation has a certain degree of rationality.
Appendix D Qualitative Results
Fig. 10 shows the qualitative results on MbgVC dataset, including generation results of V2C, Clip4Caption, SwinBert and our EAC model. V2C, Clip4Caption and SwinBert can correctly generate some actions but fail to generate correct entity names because these names never appear during training. However, with the help of the external knowledge, our EAC could well relate entities in video clips with the fine-grained actions. These cases demonstrate that our newly proposed EAC model more consistently with the requirement of practical application.

### 28 Feb 2024 v2

@@ V-D Results on Goal
TABLE VI: Ablation study on MbgVC. Ki and Kn denote the player images knowledge and player names knowledge, respectively. Ea. is the entity-aware module. Bi-GRU is bi-directional GRU module. Numbers in bold indicate the best performance.
Model Ki Kn Ea. Bi-GRU CIDEr METEOR Rouge-L BLEU-4 RoleF1
20.1 20.6 40.9 23.0 3.4
23.5 22.7 43.1 24.4 3.6
18.0 20.0 38.0 23.2 3.3
110.6 25.2 49.1 27.3 12.4
115.7 25.6 53.1 28.1 14.8
119.3 26.8 53.3 29.5 16.5
122.2 27.3 54.6 30.5 17.1
138.5 28.0 54.9 32.4 20.6
Figure 10: Data sample from the revised Goal dataset. Each video is annotated by fileid, team_id, caption, player images and player names. Each of the players involved in caption as well as their teammates serve as candidate players. The caption in this sample includes the names of players and team.
Figure 11: Qualitative results on Goal dataset. (V2C: Video2Commonsense; C4C: Clip4Caption; SwB: SwinBert; KEA: our proposed model; GT: the ground truth). The different specific entity names are marked in red. And the fine-grained actions are marked green. Since V2C does not have its own tokenizer and vocabulary list, it cannot decode names. So names are replaced with the special token $<$UNK$>$.
V-E Ablation Study
@@
To verify the contribution of the additional knowledge and other modules in KEANet, we also perform an ablation study by progressively adding these as input. The results of ablation study are shown in Table VI. Model ① consists of Resnet-18 and T5 decoder, with its input being solely video. Model ② adds the Bi-GRU module on top of Model ①. The Bi-GRU models the temporal contextual relationships between video frames, enabling the model to better understand the dynamic information in videos and predict more accurate action categories. Therefore, Model ② has shown improvements in all metrics relative to Model ①, except for the RoleF1 score. From the comparison results between of model ① and model ③, it can be observed that adding players’ images to model based on ① does not improve the model’s performance. During the training stage, if the model has not been exposed to enough specific names, or has not learned how to infer names from video and image features, it may not be able to generate these names. The comparison results between model ① and model ④ highlight a significant improvement by solely adding players’ names to the model based on ①. Providing a list of names as input allows the model to effectively integrate this information with the video features, leading to more accurate generation of names. This is likely providing the model with additional context information to help it make more accurate predictions. From the comparison results between model ④ and model ⑤, we can find that player images can emphasize the roles based on entity names and improve the names’ prediction. From the comparison results of models ⑤-⑧, it can be observed that adding entity-aware and Bi-GRU modules can improve the performance of the model. The entity-aware module can model the associations among players and focus on key players. The Bi-GRU module can model the temporal contextual information of video frame features, capturing the action information. From the above results, the additional players knowledge brings significant gains in entity awareness. The model can generate text descriptions with specific player names and fine-grained actions through additional knowledge and temporal modeling.
@@
V-F Error Analysis
The error analysis are conducted by presenting cases of KEANet on proposed VC_NBA_2022 dataset. As shown in Fig. 12, we show several errors. (1) Player Mismatching: when the number of players in the sentence is more than one, the model may decode the name of one of the players incorrectly. (2) Action Confusion: model tends to confuse similar-looking actions, such as layups and close-range two-point jump shots. (3) Lack Distance-aware Perception: model tends to confuse three-point jump shots and long-range two-point jump shots. For example, the player appears to be shooting a three-point shot, but is actually inside the three-point line.
@@ V-F Error Analysis
Given the above result, we discuss some potential ways for developing an advanced models for KEBVC task. First, beyond the current object detection, it is necessary to enhance the model’s ability to understand and handle names. Second, regarding action confusion, this may be because the model is not sensitive enough to subtle differences in the video to accurately distinguish similar actions. We need to improve the feature extraction abilities of the model to more accurately recognize similar-looking actions. Third, for the lack of distance perception, this may be because the model has trouble processing spatial information, especially when judging the relative distance of players from the basket. This indicates that we need to further improve the spatial awareness of our model.
VI Conclusion
@@
In this paper, we introduce the task of knowledge guided entity-aware basketball video captioning (KEBVC) for live text broadcast. To achieve this task, we further construct a multimodal basketball game knowledge graph (KG_NBA_2022) to provide the additional knowledge. Through partial relationships in the knowledge graph, a new multimodal basketball game video captioning dataset is then constructed. Furthermore, our experiments validate the importance of additional knowledge, including player images and entity names, for entity-aware video captioning.